# One for all: Toward unified Foundation Models for Earth vision

### v1

@@ 1 Introduction
Multiple satellites in orbit provide invaluable data essential for understanding and managing our planet. The richness of the Earth observation data lies in its diversity, encompassing various modalities such as optical, radar, multispectral, hyperspectral, and thermal imagery, each offering unique insights [1] [2]. This multiplicity is crucial in applications ranging from environmental monitoring to urban planning, demonstrating the indispensable role of remote sensing in Earth sciences.
The advent of foundation models has currently revolutionized the processing and analysis of remote sensing data [3] [4] [5] [6] [7]. Characterized by their extensive parameters and pre-trained on large-scale datasets, these models have greatly enhanced the performance on different downstream tasks. Despite their successes, current foundation models in remote sensing exhibit a critical limitation: they are predominantly tailored to either a single data modality [6] or a specific range of spatial resolutions [8] [9]. This specialization constrains their applicability across the diverse spectrum of remote sensing datasets, limiting their potential and flexibility in broader, more complex applications.
@@ 2.2 Model Design
Let $P_{s1},P_{s2},P_{naip},P_{g}$, and $P_{hyper}$ denote the patch embedding operations for five different modalities. After resizing the input images from different modalities $X_{s1}\in\mathbb{R}^{h,w,2},X_{s2}\in\mathbb{R}^{h,w,9},X_{naip}\in\mathbb{R}^{h,w,3},X_{g}\in\mathbb{R}^{h,w,4},X_{hyper}\in\mathbb{R}^{h,w,224}$ into 224x224 pixels, we can compute the embeddings simply by $E=P(X)$ for each modality.
@@
Shared Transformer Backbone. The second component is a shared Transformer backbone. In this work, this single Transformer architecture processes the embedded patches from all modalities with different spatial resolutions. This shared backbone can learn a generalized representation that is flexible and robust enough to handle not only the variety of data modalities but also their diverse spatial resolutions. In this process, we input the embeddings for each modality $E_{s1}\in\mathbb{R}^{n,d},E_{s2}\in\mathbb{R}^{n,d},E_{naip}\in\mathbb{R}^{n,d},E_{g}\in\mathbb{R}^{n,d},E_{hyper}\in\mathbb{R}^{h,w,224}$ to the Transformer backbone $F_{b}$ to learn deep features. Here $n$ represents the number of tokens and $d$ denotes the feature dimension.
@@ 3.2 Experimental Settings
Linear probing is selected as our primary evaluation technique, known for its capacity to measure the quality of representations learned by self-supervised learning. This approach involves freezing the weights of the pre-trained model and training a linear classifier on top of the representations for each task. By doing so, we can directly evaluate the discriminative power of the learned features without fully fine-tuning the entire model, thus providing insight into the model’s ability to generalize. For the classification tasks, we use top-1 Accuracy as the evaluation metric. For segmentation tasks, we use mIoU as the metric. Considering the high computational cost, for each sub-dataset, we use 10,000 data samples, with 50,000 samples in total to pre-train the OFA-Net for 100 epochs. All the experiments are conducted using PyTorch on four NVIDIA A6000 GPUs each with 48GB Memory.

### v2

@@ 1 Introduction
Multiple satellites in orbit provide invaluable data essential for understanding and managing our planet. The richness of the Earth observation data lies in its diversity, encompassing various modalities such as optical, radar, multispectral, hyperspectral, and thermal imagery, each offering unique insights [1] [2]. This multiplicity is crucial in applications ranging from environmental monitoring [3] to urban planning [4], demonstrating the indispensable role of remote sensing in Earth sciences.
The advent of foundation models has currently revolutionized the processing and analysis of remote sensing data [5, 6, 7, 8, 9, 10, 11]. Characterized by their extensive parameters and pre-trained on large-scale datasets, these models have greatly enhanced the performance on different downstream tasks. Despite their successes, current foundation models in remote sensing exhibit a critical limitation: they are predominantly tailored to either a single data modality [9] or a specific range of spatial resolutions [12] [13]. This specialization constrains their applicability across the diverse spectrum of remote sensing datasets, limiting their potential and flexibility in broader, more complex applications.
@@ 2.1 Multi-modal Dataset Construction
Let $P_{s1},P_{s2},P_{naip},P_{g}$, and $P_{hyper}$ denote the patch embedding operations for five different modalities. After resizing the input images from different modalities $X_{s1}\in\mathbb{R}^{h,w,2},X_{s2}\in\mathbb{R}^{h,w,9},X_{naip}\in\mathbb{R}^{h,w,3},X_{g}\in\mathbb{R}^{h,w,4},X_{h}\in\mathbb{R}^{h,w,224}$ into 224x224 pixels, we can compute the embeddings simply by $E=P(X)$ for each modality.
Shared Transformer Backbone. The second component is a shared Transformer backbone. In this work, this single Transformer architecture processes the embedded patches from all modalities with different spatial resolutions. This shared backbone can learn a generalized representation that is flexible and robust enough to handle not only the variety of data modalities but also their diverse spatial resolutions. In this process, we input the embeddings for each modality $E_{s1}\in\mathbb{R}^{n,d},E_{s2}\in\mathbb{R}^{n,d},E_{naip}\in\mathbb{R}^{n,d},E_{g}\in\mathbb{R}^{n,d},E_{h}\in\mathbb{R}^{h,w,224}$ to the Transformer backbone $F_{b}$ to learn deep features. Here $n$ represents the number of tokens and $d$ denotes the feature dimension.
@@ 3.2 Experimental Settings
Considering the high computational cost, we use 10,000 data samples for each sub-dataset, with 50,000 samples in total to pre-train the OFA-Net for 100 epochs. We adopt linear probing as our primary evaluation strategy, known for its capacity to measure the quality of representations learned by self-supervised learning. This approach involves freezing the weights of the pre-trained model and training a linear classifier on top of the representations for each task. By doing so, we can directly evaluate the discriminative power of the learned features without fully fine-tuning the entire model, thus providing insight into the model’s ability to generalize. A learning rate of 1e-2 is used for all the datasets. For the classification tasks, we use top-1 Accuracy as the evaluation metric. For segmentation tasks, we use mIoU as the metric. A learning rate of 1e-4 is used for all the datasets. All the experiments are conducted using PyTorch on four NVIDIA A6000 GPUs each with 48GB Memory.
@@ 4 Conclusion
5 Acknowledgement
This work is jointly supported by the German Federal Ministry of Education and Research (BMBF) in the framework of the international future AI lab “AI4EO – Artificial Intelligence for Earth Observation: Reasoning, Uncertainties, Ethics and Beyond” (grant number: 01DD20001), by German Federal Ministry for Economic Affairs and Climate Action in the framework of the “national center of excellence ML4Earth” (grant number: 50EE2201C), and by the German Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (BMUV) based on a resolution of the German Bundestag (grant number: 67KI32002B; Acronym: EKAPEx).